OpenStudy (anonymous):

[6.02] For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9

OpenStudy (kropot72):

4x - y =8 The first step to isolate y is to subtract 4x from both sides of the equation. Can you do that?

OpenStudy (kropot72):

@Mrszbieberxo Are you there?

OpenStudy (anonymous):

yes I got infinitely many solutions

OpenStudy (kropot72):

Can you please do the step that I gave previously. 4x - y =8 The first step to isolate y is to subtract 4x from both sides of the equation. Can you do that?

OpenStudy (anonymous):

y=4?

OpenStudy (kropot72):

Not really. Subtracting 4x from both sides gives: 4x - y - 4x = 8 - 4x ...............(1) Can you simplify the left hand side of (1)?

OpenStudy (anonymous):

I'm sorry i'm really confused :/

OpenStudy (kropot72):

We are looking at simplifying the left hand side of 4x - y - 4x = 8 - 4x ...............(1) If the terms on the left hand side are rearrangedas follows is the simplification clearer? (4x - 4x) - y = 8 - 4x Hint: What is 4x - 4x ?

OpenStudy (kropot72):

rearranged as*